The half-life of a radioactive tracer injected into the human body is 7 days. If a patient is given a 30-mg dose of the tracer, then how many grams of tracer are still in the patient’s body after 21 days?


1
Expert's answer
2021-03-14T19:21:38-0400

The mass of tracer are still in the patient’s body after "t = 21\space days" is:


"m = m_0 2^{-\frac{t}{T}}"

where "m_0 = 30mg" is the initial dose, "T = 7\space days" is the half-life. Thus, obtain:


"m = 30\cdot 2^{-\frac{21}{7}} = 3.75mg"

Answer. 3.75 mg.